# Radial symmetry, uniqueness and non-degeneracy of solutions to degenerate nonlinear Schrödinger equations

### v4

@@ Abstract.
Acknowledgments. T. Gou was supported by the National Natural Science Foundation of China (No. 12471113). The author would like to thank warmly the knowledgeable referee for useful remarks.
@@ 1. Introduction
Moreover, by using [10, Proposition 2], one knows that $H^{1,a}(\mathbb{R}^{d})$ is embedded compactly into $L^{q}(\mathbb{R}^{d})$ for $2<q<2_{a}^{*}$.
@@
Theorem 1.1.
([10, Theorem 1]) Let $d\geq 2$, $0<a<1$, $\omega>0$ and $2<p<2^{*}_{a}$. Then the following assertions hold true.
@@
Remark 1.1.
In fact, while $\omega\leq 0$, by using Pohozaev’s identities satisfied by solutions to (1.2) (see [10, Proposition 1]), one finds easily that there exists no solutions to (1.2) in $H^{1,a}(\mathbb{R}^{d})$.
@@
see Lemma 2.2, where $u^{*}$ denotes the symmetric-decreasing rearrangement of the function $u$. Actually, when $a=0$, then (1.5) reduces to the well-known Pólya–Szegö type inequality (see for example [1, Theorem 3.20] or [11, Lemma 7.17])
@@
where $\mu\geq 1$ is a constant, $\mathcal{H}_{d-1}$ denotes (d1)(d-1) dimensional Hausdorff measure and $u^{-1}(t):=\{x\in\mathbb{R}^{d}:u(x)=t\}$ for $t>0$. At this point, by invoking [2, Theorem 4.3], we then have that
@@
where $\chi_{A}$ denotes the characteristic function on the set $A$. It is necessary to point out that when $\mu=0$, by employing only (1.8), one cannot derive that (1.9) is valid, see for instance [3]. While $\mu\geq 1$, then the function $\tau\mapsto\tau^{\mu}$ is strictly increasing and convex on $0,+\infty)$. This jointly with (1.8) then leads to the desired result, see [2, Theorem 4.3].
@@
It is clear that $\mathcal{L}_{+,0}<\mathcal{L}_{+,1}<\mathcal{L}_{+,2}<\cdots<\mathcal{L}_{+,k}<\cdots$. As a consequence, one only needs to show that 0 is not an eigenvalue of $\mathcal{L}_{+,1}$, because Morse’s index of the ground state is $1$. For this, when $a=0$, arguing by contradiction and adapting the existing methods, one can reach a contradiction and the result follows. While $a\neq 0$, then the existing ones are not suitable to our problem and the situation becomes involved and delicate. This then forces us to discuss the non-degeneracy of the ground state in $H^{1,a}_{rad}(\mathbb{R}^{d})$.
@@ 2. Radial Symmetry
Due to $\mu\geq 1$, then the function $\tau\mapsto\tau^{\mu}$ is strictly increasing and convex on $0,+\infty)$. Moreover, since $u\in C^{\infty}(0,+\infty)$, then the boundary of $u^{-1}(t)$ is Lipschitz for any $t>0$. Invoking [2, Theorem 2.1] and (2.11), we then derive that
@@
Hence, by [2, Theorem 4.3], we have that
@@
Keeping in mind that $u$ is actually positive by the maximum principle, let us now assume by contradiction that there exists an interval $r_{1},r_{2}]\subset(0,+\infty)$ such that $u(r)=c>0$ for any $r\in[r_{1},r_{2}]$, which means that $u^{\prime\prime}(r)=0$ and $u^{\prime}(r)=0$ for any $r\in(r_{1},r_{2})$. It then infers from (2.16) that $u(r)=u^{p-1}(r)$, i.e. $u(r)=1$ for any $r\in[r_{1},r_{2}]$. Let $r_{0}\in(r_{1},r_{2})$ be such that $u(r_{0})=1$ and $u^{\prime}(r_{0})=0$. At this point, using the well-posedness of solution to the initial problem (2.16) with $u(r_{0})=1$ and $u^{\prime}(r_{0})=0$, we then have that $u(r)=1$ for any $r\geq r_{0}$. This is a contradiction, because $u$ decays exponentially at infinity by Theorem 1.1. It then yields that $u$ is strictly decreasing in the radial direction. Thus the proof is complete.
@@ 4. Non-degeneracy
Now, by taking into account [8, Lemma 9.3], we derive that there exist two linearly independent solutions $w_{1}$ and $w_{2}$ to (4.3) such that
@@
It then follows from [8, Exercise 9.6] that
@@
It then follows from [8, Theorem 9.1] that there exist two linearly independent solutions to (4.6) such that
@@
Since $v\in H_{rad}^{1,a}(\mathbb{R}^{d})$ is a nontrivial solution to (4.1), then $\mathcal{B}(v,\psi)=0$ for any $\psi\in H^{1,a}_{rad}(\mathbb{R}^{d})$. It then follows that $Tv=0$. Since $Tv=0$ is is non-oscillatory, by [12, Theorem 14.9], we then conclude that $T$ is bounded from below. Let $\varphi\in H_{rad}^{1,a}(\mathbb{R}^{d})$ be such that $(T-\lambda)\varphi=0$ for $\lambda<1$. Since $(T-\lambda)\varphi=0$ is non-oscillatory, by [12, Theorem 14.9], we then get that $\inf\sigma_{ess}(T)\geq 1$, where $\sigma_{ess}(T)$ denotes the essential spectrum of the operator $T$. Since $Tv=0$, then 0 is an eigenvalue of $T$. Moreover, we know that it is an isolated eigenvalue with finite multiplicity, because of $\sigma_{ess}(T)\geq 1$. Observe that
@@
Hence we know that $v$ changes sign in $(0,+\infty)$, because of $u>0$. It then follows from [12, Theorem 14.10] that $v$ is not an eigenfunction of the smallest eigenvalue of $T$. As a consequence, we derive that $T$ admits at least one negative eigenvalue. Indeed, we are able to show that there exists only one negative eigenvalue for the operator $T$. Contrarily, we assume that there exist two negative eigenvalues $\lambda_{2}<\lambda_{1}<0$ corresponding to the operator $T$. Therefore, there exist nontrivial $\varphi_{1},\varphi_{2}\in H_{rad}^{1,a}(\mathbb{R}^{d})$ such that
@@
which is impossible. Thereby, we conclude that $T$ has only one negative eigenvalue. It then infers that 0 is the second eigenvalue of the operator $T$. At this point, using [12, Theorem 14.10], we then derive that $v$ has exactly one zero in $(0,+\infty)$. This completes the proof. ∎

### v5

@@ Abstract.
Acknowledgments. T. Gou was supported by the National Natural Science Foundation of China (No. 12471113). The author warmly thanks Profs. Rupert Frank and Jean Dolbeault for useful discussions and for bringing the literature [1, 9] to his attention. He also thanks Prof. Rupert Frank for pointing out a mistake presented in the original proof of Theorem 2.1. The author also would like to thank warmly the knowledgeable referee for useful remarks.
@@ 1. Introduction
Moreover, by using [12, Proposition 2], one knows that $H^{1,a}(\mathbb{R}^{d})$ is embedded compactly into $L^{q}(\mathbb{R}^{d})$ for $2<q<2_{a}^{*}$.
@@
Theorem $1.$
([12, Theorem 1]) Let $d\geq 2$, $0<a<1$, $\omega>0$ and $2<p<2^{*}_{a}$. Then the following assertions hold true.
@@
Remark $1.$
In fact, while $\omega\leq 0$, by using Pohozaev’s identities satisfied by solutions to (1.2) (see [12, Proposition 1]), one finds easily that there exists no solutions to (1.2) in $H^{1,a}(\mathbb{R}^{d})$.
@@
see Lemma 2.2, where $u^{*}$ denotes the symmetric-decreasing rearrangement of the function $u$. Actually, when $a=0$, then (1.5) reduces to the well-known Pólya–Szegö type inequality (see for example [2, Theorem 3.20] or [13, Lemma 7.17])
@@
where $\mu\geq 1$ is a constant, $\mathcal{H}_{d-1}$ denotes (d1)(d-1) dimensional Hausdorff measure and $u^{-1}(t):=\{x\in\mathbb{R}^{d}:u(x)=t\}$ for $t>0$. At this point, by invoking [3, Theorem 4.3], we then have that
@@
where $\chi_{A}$ denotes the characteristic function on the set AA. It is necessary to point out that when $\mu=0$, by employing only (1.8), one cannot derive that (1.9) is valid, see for instance [4]. While $\mu\geq 1$, then the function $\tau\mapsto\tau^{\mu}$ is strictly increasing and convex on $0,+\infty)$. This jointly with (1.8) then leads to the desired result, see [3, Theorem 4.3].
@@
It is clear that $\mathcal{L}_{+,0}<\mathcal{L}_{+,1}<\mathcal{L}_{+,2}<\cdots<\mathcal{L}_{+,k}<\cdots$. As a consequence, one only needs to show that 0 is not an eigenvalue of $\mathcal{L}_{+,1}$, because Morse’s index of the ground state is 11. For this, when $a=0$, arguing by contradiction and adapting the existing methods, one can reach a contradiction and the result follows. While $a\neq 0$, then the existing ones are not suitable to our problem and the situation becomes involved and delicate. This then forces us to discuss the non-degeneracy of the ground state in $H^{1,a}_{rad}(\mathbb{R}^{d})$.
@@ 2. Radial Symmetry
Due to $\mu\geq 1$, then the function $\tau\mapsto\tau^{\mu}$ is strictly increasing and convex on $0,+\infty)$. Moreover, since $u\in C^{\infty}(0,+\infty)$, then the boundary of $u^{-1}(t)$ is Lipschitz for any $t>0$. Invoking [3, Theorem 2.1] and (2.11), we then derive that
@@
Hence, by [3, Theorem 4.3], we have that
@@
Keeping in mind that $u$ is actually positive by the maximum principle, let us now assume by contradiction that there exists an interval $r_{1},r_{2}]\subset(0,+\infty)$ such that $u(r)=c>0$ for any $r\in[r_{1},r_{2}]$, which means that $u^{\prime\prime}(r)=0$ and $u^{\prime}(r)=0$ for any $r\in(r_{1},r_{2})$. It then infers from (2.16) that $u(r)=u^{p-1}(r)$, i.e. $u(r)=1$ for any $r\in[r_{1},r_{2}]$. Let $r_{0}\in(r_{1},r_{2})$ be such that $u(r_{0})=1$ and $u^{\prime}(r_{0})=0$. At this point, using the well-posedness of solution to the initial problem (2.16) with $u(r_{0})=1$ and $u^{\prime}(r_{0})=0$, we then have that $u(r)=1$ for any $r\geq r_{0}$. This is a contradiction, because $u$ decays exponentially at infinity by Theorem $1.$ It then yields that $u$ is strictly decreasing in the radial direction. Thus the proof is complete.
@@ 4. Non-degeneracy
Now, by taking into account [10, Lemma 9.3], we derive that there exist two linearly independent solutions $w_{1}$ and $w_{2}$ to (4.3) such that
@@
It then follows from [10, Exercise 9.6] that
@@
It then follows from [10, Theorem 9.1] that there exist two linearly independent solutions to (4.6) such that
@@
Since $v\in H_{rad}^{1,a}(\mathbb{R}^{d})$ is a nontrivial solution to (4.1), then $\mathcal{B}(v,\psi)=0$ for any $\psi\in H^{1,a}_{rad}(\mathbb{R}^{d})$. It then follows that $Tv=0$. Since $Tv=0$ is is non-oscillatory, by [14, Theorem 14.9], we then conclude that $T$ is bounded from below. Let $\varphi\in H_{rad}^{1,a}(\mathbb{R}^{d})$ be such that $(T-\lambda)\varphi=0$ for $\lambda<1$. Since $(T-\lambda)\varphi=0$ is non-oscillatory, by [14, Theorem 14.9], we then get that $\inf\sigma_{ess}(T)\geq 1$, where $\sigma_{ess}(T)$ denotes the essential spectrum of the operator $T$. Since $Tv=0$, then 0 is an eigenvalue of $T$. Moreover, we know that it is an isolated eigenvalue with finite multiplicity, because of $\sigma_{ess}(T)\geq 1$. Observe that
@@
Hence we know that $v$ changes sign in $(0,+\infty)$, because of $u>0$. It then follows from [14, Theorem 14.10] that $v$ is not an eigenfunction of the smallest eigenvalue of $T$. As a consequence, we derive that $T$ admits at least one negative eigenvalue. Indeed, we are able to show that there exists only one negative eigenvalue for the operator $T$. Contrarily, we assume that there exist two negative eigenvalues $\lambda_{2}<\lambda_{1}<0$ corresponding to the operator $T$. Therefore, there exist nontrivial $\varphi_{1},\varphi_{2}\in H_{rad}^{1,a}(\mathbb{R}^{d})$ such that
@@
which is impossible. Thereby, we conclude that $T$ has only one negative eigenvalue. It then infers that 0 is the second eigenvalue of the operator $T$. At this point, using [14, Theorem 14.10], we then derive that $v$ has exactly one zero in $(0,+\infty)$. This completes the proof. ∎